What weight of lactic acid must be dissolved in a liter of solution to produce a hydronium-ion concentration of 1×10^-5


1
Expert's answer
2022-03-16T18:51:04-0400

Solution:

lactic acid = CH3CH(OH)COOH

hydronium ion = H3O+


Lactic acid (CH3CH(OH)COOH) is a weak acid because it partially dissociates in an aqueous solution or it does not ionize completely in solution to produce hydronium ions (H3O+).

Equation representing the dissociation of CH3CH(OH)COOH:

CH3CH(OH)COOH(aq) + H2O(l) ⇋ H3O+(aq) + CH3CH(OH)COO(aq)


CH3CH(OH)COOH = HA

CH3CH(OH)COO = A


ICE Table:



According to the ICE Table:

[H3O+] = [CH3CH(OH)COO] = 1×10−5

[CH3CH(OH)COOH] = C − 1×10−5


The Ka expression for the dissociation of CH3CH(OH)COOH:



Ka = 1.4×10−4 (for lactic acid )

Therefore,

1.4×10−4 = [1×10−5] × [1×10−5] / [C − 1×10−5]

7.14×10−7 = C − 1×10−5

C = 1.07×10−5


Molarity of lactic acid = 1.07×10−5 M

Moles of lactic acid = Molarity of lactic acid × Liters of solution

Therefore,

Moles of lactic acid = (1.07×10−5 M) × (1 L) = 1.07×10−5 mol


The molar mass of lactic acid (CH3CH(OH)COOH) is 90.08 g/mol

Therefore,

Mass of lactic acid = (1.07×10−5 mol) × (90.08 g/mol) = 9.64×10−4 g

Mass of lactic acid = 9.64×10−4 g

Answer: 9.64×10−4 grams of lactic acid must be dissolved